# Linying Lv

### v3

@@ B. Additional Tables
Panel A: Firm Level Measures
@@
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
@@
Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
@@
Panel B: Industry Level Measures
@@
Time trend (ttr) The number of years elapsed from the beginning of the sample.
Panel D: Report Level Measures
$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
@@
This table reports summary statistics of numerical measures.
Mean Std p25 P50 P75 N
@@
This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed. The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both. The estimation targets are the three-day $CAR$ centered around the analyst report release date. The machine learning model employed is Ridge regression, with training samples expanding yearly. The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al. (2020). In columns (2), (4), and (6), the benchmark estimation is set to zero. Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
@@
This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows. The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year. $CAR$ measures are centered on the release day of analyst reports ($T_{0}$). Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
@@
Panel A: Full sample period
@@
Panel B: Earnings announcement period
@@
Panel C: Non-earnings announcement period
@@
This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
year Shops Other Manuf Chems
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
year Durbl BusEq Hlth NoDur
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
year Money Telcm Enrgy Utils
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars. All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI). The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4. Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
@@ D. Theory
A Single Auction Model
@@
The profit of the insider can be written as:
@@
Solving for optimal profit, I get:
@@
Hence, I can express $\alpha$ and $\beta$ :
@@
I now look at the price-setting rule:
@@
Hence, I can express $\lambda$ and $\mu$ :
@@ E. Delta Method Approximation
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
@@
Substituting the partial derivatives into the Taylor expansion, I obtain:
@@
Taking the expectation on both sides:
@@
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
@@
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
@@
Substituting the partial derivatives:
@@
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as:
@@
and the mean price impact per dollar over subsample $s$ as:
@@
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
@@
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
@@
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:

### v4

@@ B. Additional Tables
Panel A: Firm-Level Measures
@@
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of the quarter prior to the report release.
PriorCAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within a 90-day window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
@@
DD The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
@@
Panel B: Industry-Level Measures
@@
TimeTrend The number of years elapsed from the beginning of the sample.
Panel D: Report-Level Measures
$REC_{REV}$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF_{REV}$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP_{REV}$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
@@
This table reports summary statistics of numerical measures. See Table A2 for detailed variable definitions.
Mean Std. Dev. p25 P50 P75 N
@@
This table presents an analysis of analyst reports where all numerical values have been removed from the text, designed to isolate the contribution of pure qualitative information. I report the $R^{2}_{\mathrm{OOS}}$ from Ridge regressions predicting three-day cumulative abnormal returns ($CAR[-1,+1]$) using several input configurations: numerical forecast revisions only, embeddings from the number-free text, and a combination of both. The table also includes pairwise tests to assess whether the number-free text contains more information than the numerical revisions alone (Column (3)-(1)), and whether those revisions provide incremental value when added to the number-free text (Column (5)-(3)). T-statistics for the $R^{2}_{\mathrm{OOS}}$ and their differences are calculated following the procedure outlined by Gu et al. (2020).
Year Rev only t-stat Text only t-stat Rev + text t-stat t-stat t-stat
@@
Table A5: Information Content of Analyst Reports using Alternative Machine Learning Models
This table compares the performance ($R^{2}_{\mathrm{OOS}}$) of several alternative machine learning models in estimating the information content of analyst report text. The models tested include Partial Least Squares (PLS), Extreme Gradient Boosting (XGBoost), and a series of Neural Networks (NN1-NN5) with an increasing number of hidden layers. The out-of-sample $R^{2}$ for each model is calculated annually using an expanding training window that begins in 2000. The ‘Overall’ row summarizes the performance across the full 2015–2023 sample period. T-statistics for the $R^{2}_{\mathrm{OOS}}$ are calculated against a zero benchmark following the procedure in Gu et al. (2020).
PLS XGBoost NN1 NN2 NN3 NN4 NN5
year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
2015 12.23% 10.95 8.87% 11.25 15.81% 9.72 13.84% 14.53 15.30% 14.10 11.58% 12.32 16.61% 16.95
2016 10.27% 15.59 6.66% 5.63 12.59% 24.64 11.78% 13.02 10.99% 9.10 12.03% 14.75 12.46% 17.41
2017 6.60% 6.94 4.92% 5.26 9.08% 4.64 10.01% 4.64 9.58% 4.76 8.57% 4.51 10.03% 5.49
2018 8.08% 13.37 5.15% 8.07 10.73% 26.65 10.54% 16.42 10.41% 11.43 9.86% 17.44 11.98% 13.97
2019 11.54% 38.70 6.32% 16.40 13.43% 17.37 15.18% 14.73 15.59% 18.06 15.89% 10.48 15.37% 16.94
2020 3.21% 2.16 3.02% 3.64 4.45% 2.92 4.73% 2.78 3.85% 2.27 4.25% 3.08 5.51% 4.06
2021 2.14% 1.76 4.30% 3.34 4.07% 6.44 5.88% 7.13 9.28% 13.59 8.47% 9.44 8.83% 7.67
2022 11.42% 7.40 10.04% 7.80 13.78% 8.31 15.67% 6.38 16.44% 6.64 17.63% 8.55 17.56% 7.12
2023 8.14% 9.26 4.23% 3.82 9.54% 8.50 12.16% 10.76 11.06% 8.59 12.18% 11.59 11.70% 6.78
Overall 8.26% 6.66 5.97% 6.76 10.42% 8.11 11.04% 7.54 11.18% 7.29 11.05% 6.62 12.12% 7.93
@@
This table reports the $R^{2}_{\mathrm{OOS}}$ from Ridge regression models that use analyst report text embeddings to estimate CAR across alternative event windows. The target variables are abnormal returns centered on the report release date ($T_{0}$): same-day abnormal return (AR[0]), two-day cumulative return (CAR[0,+1]), and five-day cumulative return (CAR[–2,+2]). All $R^{2}_{\mathrm{OOS}}$ values are computed annually using an expanding training sample starting in 2000. Panel A shows results for the full sample, while Panels B and C restrict the sample to reports released within one day (excluding same day) of an earnings announcement and those outside this window, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
@@
Panel A: Full sample
@@
Panel B: Reports issued within earnings announcement window
@@
Panel C: Reports issued beyond earnings announcement window
@@
This table reports information content of analyst report text for each of the Fama-French 12 (FF12) industries for the period of 2015–2023. Industry definitions are based on Kenneth French’s data library. T-statistics for all $R^{2}_{\mathrm{OOS}}$ values are calculated against a zero benchmark following the procedure in Gu et al. (2020).
Shops Other Manuf Chems
year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
Durbl BusEq Hlth NoDur
year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
Money Telcm Enrgy Utils
year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
@@
Table A8: Summary Statistics of Alternative Analyst Information Value Estimations
This table reports summary statistics for several alternative estimations of analyst information value for S&P 100 stocks from 2015Q1 to 2023Q4. These alternatives modify different components of the main information value calculation in Table 8 (defined as explained return variance divided by price impact). The ‘TAQ volatility’ specification calculates explained variance as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$ using one-minute log returns, while the ‘Intraday’ version excludes overnight returns from this calculation. The ‘EMO’ and ‘CLNV’ specifications use alternative trade-signing algorithms from Ellis et al. (2000) and Lee and Ready (1991) to calculate price impact. For each specification, the table reports the mean, standard error (SE), and 95% and 99% confidence intervals, with all estimates computed with the delta method and expressed in millions of 2020 dollars.
@@ D. Theory
Single Auction Model
@@
The profit of the insider can be written as
@@
Solving for optimal profit, I get
@@
Hence, I can express $\alpha$ and $\beta$ as
@@
I now look at the price-setting rule as
@@
Hence, I can express $\lambda$ and $\mu$ as
@@ E. Delta Method Approximation
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are
@@
Substituting the partial derivatives into the Taylor expansion, I obtain
@@
Taking expectations on both sides gives
@@
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation can be simplified to
@@
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by
@@
Substituting the partial derivatives yields
@@
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as
@@
and the mean price impact per dollar over subsample $s$ as
@@
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by
@@
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as
@@
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as